OpenStudy (anonymous):

Please explain how to convert these percents to a fraction. 33.3% 83.3% 62.5% 87.5%

OpenStudy (jack1):

percent means per 100 so 33.3 % = 33.3/100 can be simplified to 3.33/10, 0.33/1...etc

OpenStudy (anonymous):

But changing a decimal to a fraction is confusing me.

OpenStudy (anonymous):

move the decimal point to the left two places

OpenStudy (goldphenoix):

It's easy. Decimal means out of 100.

OpenStudy (goldphenoix):

So any decimal, out of 100.

OpenStudy (anonymous):

then turn the decimal in to a fraction if you need to do that

OpenStudy (anonymous):

Okay let me try that

OpenStudy (jack1):

"Decimal means out of 100" ... dont u mean percent...?

OpenStudy (goldphenoix):

^My bad.

OpenStudy (anonymous):

How do I figure out the denominator?

OpenStudy (whpalmer4):

Okay, it's helpful to know the decimal equivalents of some common fractions: 1/2 = 0.5 1/3 = 0.33333... 1/4 = 0.25 1/5 = 0.2 1/6 = 0.16666... 1/7 = 0.142857142857... 1/8 = 0.125 1/9 = 0.1111111... 1/10 = 0.1 1/16 = 0.0625 1/20 = 0.05 1/25 = 0.04 1/32 = 0.03125 2/3 = 2*1/3 = 0.666666... 2/5 = 2*1/5 = 0.4 3/5 = 3*1/5 = 0.6 5/6 = 5*1/6 = 0.83333... 2/7 = 0.285714... etc. So looking at your list, 62.5% is going to be some multiple of 1/16 = 0.0625 87.5% is a multiple of 0.125 33.3% should be obvious 83.3% is a multiple of 0.1666... you can also just try multiplying the decimal by increasingly large integers on your calculator until you find one that makes the result "come out even" 3*0.333333 = 0.99999 -> that's exceedingly close to 1, so 1/3 would be the fraction.